OpenStudy (anonymous):

Please help! Find an equation in slope-intercept form for the line through points (-1, -4) and (3, 2)

OpenStudy (karatechopper):

Do you know the equation for Slope Intercept Form?

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

Don't i need to put it in point slope first?

OpenStudy (karatechopper):

Good! So..Do you know how to find the slope of a line?

OpenStudy (karatechopper):

U don't have to. I can show u an easier way...if you want.

OpenStudy (anonymous):

I know the slope formula, should i do that?

OpenStudy (anonymous):

\[M = \frac{y2-y1}{x2-x1} \] This is the gradient formula, and you can use this when there are two points and in this case there are two points provided of (-1,-4) and (3,2).

OpenStudy (anonymous):

(-1,-4) and (3,2) (x1,y1) (x2,y2) now try using these values and substituting them into the gradient formula above.

OpenStudy (anonymous):

So slope = 3/2

OpenStudy (anonymous):

Does it matter which ordered pair i plug into y=mx+b?

OpenStudy (karatechopper):

Oh..ok i guess msamido can take it up for me:)

OpenStudy (anonymous):

Well I have my slope, all i need now is the ordered pair to plug into the final equation right?

OpenStudy (anonymous):

(-1, -4) and (3, 2) U recall the equation for gradient y2-y1=m(x2-x1) (2+1)=m(3+4) m=3/7 then take equation y-y1=m(x-x1) y-2=(7/3)(x-3) y=((7x)/3)-5

OpenStudy (anonymous):

But why did I go the slope formula and not get the same slope as you? Shouldn't both ways work?

OpenStudy (anonymous):

post what U got @smagyar

OpenStudy (anonymous):

\[(2+4) / (3+1) = 3/2\]

OpenStudy (anonymous):

i had seen my mistake....i had exchanged da variabls sumwhr...but tried 2 correct it but it ddnt post ...i wll fix it

OpenStudy (anonymous):

(-1, -4) and (3, 2) U recall the equation for gradient y2-y1=m(x2-x1) (2+4)=m(3+1) m=3/2 then take equation y-y1=m(x-x1) y-2=(3/2)(x-3) y=((3x)/2)-(5/2)

OpenStudy (anonymous):

i hope is now clear

OpenStudy (anonymous):

Thank you!!